Smith Distributors, Inc., supplies ice cream shops with various toppings for making sundaes. On November 17, 2021, a fire resulted in the loss of all of the toppings stored in one section of the warehouse. The company must provide its insurance company with an estimate of the amount of inventory lost. The following information is available from the company's accounting records:

Fruit
Toppings
Marshmallow
Toppings
Chocolate
Toppings
Inventory, January 1, 2021 $ 29,000 $ 7,900 $ 3,900
Net purchases through Nov. 17 195,000 45,000 12,900
Net sales through Nov. 17 245,000 64,000 20,900
Historical gross profit ratio 15 % 25 % 30 %
Toppings Estimated Loss of inventory
Fruit
Marshmallow
Chocolate
Answer #1

1) calculate estimated cost of lost in the fire

Fruit Marshmallow Chocolate
Beginning inventory 29000 7900 3900
Purchase 195000 45000 12900
Cost of goods available for sale 224000 52900 16800
Less: Cost of goods sold 245000*85% = -208250 64000*75% = -48000 20900*70% = -14630
Inventory lost in fire 15750 4900 2170